OpenStudy (marcoreus11):

Help! I don't know what to do please show me a example for each of them please

OpenStudy (marcoreus11):

Help

OpenStudy (marcoreus11):

@jim_thompson5910 can u please help me again if u dont mind? can u show me an example for each, i am pretty sure after that i will get the point how to do the others thx

jimthompson5910 (jim_thompson5910):

Problem 13 "3 units up" means that the y coordinate will change. The x coordinate of each point stays the same. Simply add 3 to each y value in the table to get the new updated table. the first y value of 3 turns into 3+3 = 6 the second y value of 1 turns into 1+3 = 4 the third y value of -2 turns into -2+3 = 1 the fourth y value of -1 turns into -1+3 = 2

jimthompson5910 (jim_thompson5910):

problems 14 and 15 are similar. For 14, you'll subtract 1 from each y value. For 15 you'll add 4 to each y value

OpenStudy (marcoreus11):

thank you what about 16? @jim_thompson5910

OpenStudy (mathmale):

Can you apply any of what you've just seen / learned to attacking #16?

OpenStudy (mathmale):

The approach you need to take is exactly the same, other than that you have to deal with fractions in #16.

OpenStudy (marcoreus11):

@mathmale can i use what ever number i want for y? then subtract it?

OpenStudy (mathmale):

No. Look at the four given f(x) values, and then subtract 2/3 from each one, to form a new table.

OpenStudy (marcoreus11):

they dont give u a table

OpenStudy (mathmale):

On the contrary! You shared a table that was given to you at the beginning.

OpenStudy (marcoreus11):

where?

OpenStudy (marcoreus11):

i am so lost

OpenStudy (mathmale):

http://assets.openstudy.com/updates/attachments/57e46b1be4b0dc03c974babd-marcoreus11-1474587429476-image.jpeg You provided this illustration yourself.

OpenStudy (marcoreus11):

yeah i know but those pictures are for #13 14 and 15

OpenStudy (marcoreus11):

we are doing 16

jimthompson5910 (jim_thompson5910):

@MarcoReus11 is there something below #16 that is cut off?

OpenStudy (mathmale):

You'll need to be more specific about what you want help with. You didn't tell me you were moving on to a new problem.

OpenStudy (marcoreus11):

no @jim_thompson5910

OpenStudy (mathmale):

If you begin with y=f(x) and the translation is "2/3 unit down," then the new function is represented by y=f(x)-(2/3).

OpenStudy (mathmale):

Would you try #17 now, please?

OpenStudy (marcoreus11):

y=f(x)+(4)

OpenStudy (marcoreus11):

and my bad we are making a equation i misread the direction thought it said make a table

OpenStudy (marcoreus11):

y=f(x)+2

OpenStudy (mathmale):

that's fine. You don't actually need the parentheses, but the parentheses don't hurt.

OpenStudy (mathmale):

Yes, your most recent input is correct for #18.

OpenStudy (mathmale):

Any further questions about this problem?

OpenStudy (marcoreus11):

no and thank you so much and showing me what i missed

OpenStudy (mathmale):

My pleasure!

OpenStudy (marcoreus11):

wait i have another problem that says translate 2 units left, 7 unit down can u check if i did this right using y=f(x)

OpenStudy (marcoreus11):

y=f(-2)-7

OpenStudy (mathmale):

Your -7 for "7 units down" is fine. Regarding a horiz. translation to the left, try this: y=f(x+2)-7 You cannot disregard or drop the x.

OpenStudy (marcoreus11):

oh okay but if it said 2 units right it would be y=f(x-2)-7 right?

OpenStudy (marcoreus11):

just making sure cause i heard a teacher mentioning something about it being opposite

OpenStudy (mathmale):

Yes, you're correct. If you begin with y=f(x) and wish to translate the graph 2 units to the right and 7 units down, then you're looking at y=f(x-2)-7.

OpenStudy (marcoreus11):

how would i reflect it across the x axis?

OpenStudy (marcoreus11):

translated 5 units right, reflected across the x axis

OpenStudy (marcoreus11):

i understand the translate 5 units right

OpenStudy (mathmale):

If you begin with y=f(x), putting a neg sign in front of f(x) will do that. If you begin with y=f(x) and translate the graph 5 units to the right, then you get y=f(x-5). Writing the "-" symbol in front of that will reflect the shifted graph in the x-axis.

OpenStudy (mathmale):

Remember, you can check your work by actually writing out specific formulas. Graph y=f(x)=x^2. Then graph y=f(x-5)=(x-5)^2. Then graph y=-f(x-5).

OpenStudy (marcoreus11):

i am just trying to make sure i know everything so if i were to start off with y=-f(x) and it said translate it to the x axis i would just remove the -(neg) symbol right?

OpenStudy (mathmale):